40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5355     Entry time: Wed Sep 7 08:14:01 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: final OSEM check 

All fine, except ITMX_sensor_UL's  60  counts deep hoop for an hour.

Attachment 1: finalcheck.jpg  116 kB  | Hide | Hide all
finalcheck.jpg
Attachment 2: ITMX10min.jpg  68 kB  Uploaded Wed Sep 7 10:08:05 2011  | Hide | Hide all
ITMX10min.jpg
Attachment 3: finalsum.png  41 kB  Uploaded Wed Sep 7 13:07:26 2011  | Hide | Hide all
finalsum.png
ELOG V3.1.3-